An unstable particle at rest breaks up into two fragments of unequal mass. The mass of the lighter fragment is equal to 2.30 10-28 kg and that of the heavier fragment is 1.65 10-27 kg. If the lighter fragment has a speed of 0.893c after the breakup, what is the speed of the heavier fragment?
Answer #1

The force exerted on each fragment is the same, what you are dealing with is the momentum of each object. Since the force exerted on each object is equal, the impulse (Ns) is equal for each piece and Ns1 = Ns2 <==> mv1 = mv2

(2.30 x 10^28)(0.893c) = (1.65 x 10^27)(v2)
v2 = (2.30 x 10^28)(0.893c) / (1.60 x 10^27)
v2 = 12.83c
Answer #2

2.3*10^-28*0.893c=1.65*10^-27*v

v=0.1245c
Answer #3

The force exerted on each fragment is the same, what you are dealing with is the momentum of each object. Since the force exerted on each object is equal, the impulse (Ns) is equal for each piece and Ns1 = Ns2 <==> mv1 = mv2

(2.30 x 10^28)(0.893c) = (1.60 x 10^27)(v2)
v2 = (2.30 x 10^28)(0.893c) / (1.60 x 10^27)
v2 = 12.84c